# Ten (or more!) reasons to register your software with the Astrophysics Source Code Library

### v1

@@ Abstract
This presentation covered the benefits of registering astronomy research software with the Astrophysics Source Code Library (ASCL, ascl.net), a free online registry for software used in astronomy research. Indexed by ADS and Clarivate’s Web of Science, the ASCL currently contains over 3600 codes, and its entries have been cited over 17,000 times. Registering your code with the ASCL is easy with our online submissions system. Making your software available for examination shows confidence in your research and makes your research more transparent, reproducible, and falsifiable. ASCL registration allows your software to be cited on its own merits and provides a citation method that is trackable and accepted by all astronomy journals, and by journals such as Science and Nature. Adding your code to the ASCL also allows others to find your code more easily, as it can then be found not only in the ASCL itself, but also in ADS, Web of Science, and Google Scholar.
@@ 1 Introduction
The Astrophysics Source Code Library [^1] https://ascl.net (ASCL) is a free online repository and registry of source codes used in astronomy research that have appeared in or been submitted to peer-reviewed publications. Established at Michigan Technological University in 1999 by Robert Nemiroff and John Wallin (Nemiroff & Wallin 1999), its editors examine peer-reviewed papers for computational methods and create entries for the found software to the library. The ASCL also accepts submitted entries that meet its criteria. [^2] https://ascl.net/home/getwp/3593 With over 3,500 entries, the ASCL covers a significant number of the astrophysics source codes used in peer-reviewed studies.
@@ 2.1 Get a unique identifier: an ASCL ID
An ASCL ID is assigned when a code is accepted by the ASCL. This unique identifier is recognized by NASA’s Astrophysics Data System (ADS), the resolution service Identifiers.org, [^3] https://registry.identifiers.org/registry/ascl the Web of Science Data Citation Index, [^4] https://clarivate.com/academia-government/scientific-and-academic-research/research-discovery-and-referencing/web-of-science/data-citation-index/ Science, Nature, and other journals, and by publishers including AAS Publications and Elsevier.
@@ 2.3 Makes your code citable on its own merit
Because the ASCL ID is unique to a code, it can be, and often is, used to cite computational methods in journal articles, proceedings, and in theses. The ability to cite software is just as important as the ability to cite a scientific paper; the software citation provides credit to the author(s) of the code and improves the transparency and reproducibility of the research for which it was used.
When a journal article has been written about a piece of software, that software paper is often used to cite the code. However, the FORCE11 Software Citation Guidelines (Smith et al. 2016), which many publishers support, state that “Software citations should be accorded the same importance in the scholarly record as citations of other research products, such as publications and data”. The FORCE11 guidelines also recognize the value of software papers, and say that “a request from the software authors to cite a paper should typically be respected, and the paper cited in addition to the software.” (Emphasis ours.) The ASCL strongly supports this, and when a code author has indicated how their code should be cited, that specification is included in ASCL entries in the “Preferred citation method” field.
Many codes do not have a dedicated software paper, instead having been written to support research. Although the code author will sometimes request that the research paper be cited if the software is used, more often than not there is no clear way to cite the code. The ASCL can help with that, as the ASCL ID can be used for the citation. [^5] https://ascl.net/wordpress/about-ascl/citing-ascl-code-entries/
@@ 2.11 Makes astronomy research more falsifiable, transparent, reproducible … better!
Science is a “show me” endeavor. Computational methods are methods, and science demands that research methods be revealed. This allows others not only to see how your research was conducted, but also enables others to verify your methods and build upon them.

### v2

@@ Abstract
This presentation covered the benefits of registering astronomy research software with the Astrophysics Source Code Library (ASCL, ascl.net), a free online registry for software used in astronomy research. Indexed by ADS and Clarivate’s Web of Science, the ASCL currently contains over 3,500 codes, and its entries have been cited over 17,000 times. Registering your code with the ASCL is easy with our online submissions system. Making your software available for examination shows confidence in your research and makes your research more transparent, reproducible, and falsifiable. ASCL registration allows your software to be cited on its own merits and provides a citation method that is trackable and accepted by all astronomy journals, and by journals such as Science and Nature. Adding your code to the ASCL also allows others to find your code more easily, as it can then be found not only in the ASCL itself, but also in ADS, Web of Science, and Google Scholar.
@@ 1 Introduction
The Astrophysics Source Code Library [^1] https://ascl.net (ASCL) is a free online repository and registry of source codes used in astronomy research that have appeared in or been submitted to peer-reviewed publications. Established at Michigan Technological University in 1999 by Robert Nemiroff and John Wallin (Nemiroff & Wallin 1999), its editors examine peer-reviewed papers for computational methods and create entries for the found software to the library. The ASCL also accepts submitted entries that meet its criteria. [^2] https://ascl.net/home/getwp/3593 With over 3,500 entries, it covers a significant number of the astrophysics source codes used in peer-reviewed studies.
@@ 2.1 Get a unique identifier: an ASCL ID
An ASCL ID is assigned after an entry is vetted and accepted by the ASCL. This unique identifier is recognized by NASA’s Astrophysics Data System (ADS), the resolution service Identifiers.org, [^3] https://registry.identifiers.org/registry/ascl the Web of Science Data Citation Index, [^4] https://clarivate.com/academia-government/scientific-and-academic-research/research-discovery-and-referencing/web-of-science/data-citation-index/ Science, Nature, and other journals, and by publishers including AAS Publications and Elsevier.
@@ 2.3 Makes your code citable on its own merit
Because the ASCL ID is unique to a code, it can be, and often is, used to cite computational methods in journal articles, proceedings, and theses. The ability to cite software is just as important as the ability to cite a scientific paper; the software citation provides credit to the author(s) of the code and improves the transparency and reproducibility of the research for which it was used.
When a journal article has been written about a piece of software, that software paper is often used to cite the code. However, the FORCE11 Software Citation Guidelines (Smith et al. 2016), which many publishers support, state that “Software citations should be accorded the same importance in the scholarly record as citations of other research products, such as publications and data”. The FORCE11 guidelines also recognize the value of software papers, and say that “a request from the software authors to cite a paper should typically be respected, and the paper cited in addition to the software.” (Emphasis ours.) The ASCL strongly supports this; when a code author has indicated how their code should be cited, that specification is shown in the “Preferred citation method” field of the ASCL entry for that software.
Many codes do not have a dedicated software paper, instead having been written to support research. Although the code author will sometimes request that the research paper be cited if the software is used, there is often no clear way to cite the code. The ASCL can help with that, as the ASCL ID can be used for the citation. [^5] https://ascl.net/wordpress/about-ascl/citing-ascl-code-entries/
@@ 2.11 Makes astronomy research more falsifiable, transparent, reproducible … better!
Science is a “show me” endeavor. Computational methods are methods, and science demands that research methods be revealed. This allows others not only to see how your research was conducted, but also enables others to verify your work and build upon it.